A traain is crossing an observer standing on platform. The first compartment of the train takes 2 second to cross the observer while the second compartment takes 2.5 seconds to cross him. The train is moving with uniform acceleration. Find: (a) the velocity of the train when the front of the first compartment crosses the observer, and (b) the acceleration of the train. Given, the length of each compartment is 15 metres.

Grade:11

1 Answers

Mohit Chhabra
8 years ago
Let speed of train is u at the front of first compartment and v at the end of first and front of second compartment
let a be the acceleration
as 1 compartment takes 2 sec to cross the observer apply 1 eq of motion
v=u+2a [eq 1]
apply 2 eq of motion
15=2u+2a [eq 2]
 
now 2 compart takes 2.5 sct to cross the observer
apply 2 eq of motion
15=2.5v+3.125a  [eq 3]
 
solving the above 3 eq we get
u=147/18  a= -6/9